FUNDAMENTAL AND TORSIONAL COMBINATION BANDS OF N_{2}O-C_{2}H_{2} AND N_{2}O-C_{2}D_{2} IN THE N_{2}O ν_{1} REGION

\underline{M. DEHGHANY}, MAHIN AFSHARI, J. N. OLIAEE, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, Calgary, AB T2N 1N4, CANADA; A.R.W. MCKEL-LAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON K1A 0R6, CANADA.

Spectra of the weakly-bound N_{2}O-C_{2}H_{2} and N_{2}O-C_{2}D_{2} complexes in the region of the N_{2}O ν_{1} fundamental band ( 2224 cm^{−1}) are observed in a pulsed supersonic slit jet expansion probed with a tunable diode laser. Two bands are analyzed for each complex: the fundamental (N-N stretch), and a combination involving the intermolecular torsional (out-of-plane bend) vibration. The resulting torsional frequencies are 44.37 and 40.01 cm^{−1} for the C_{2}H_{2} and C_{2}D_{2} complexes, respectively. This represents the first observation of the N_{2}O-C_{2}D_{2} isotopomer, and the first direct determination of an intermolecular frequency for nitrous oxide - acetylene.